### v1

# Discover the GLM theory on four pages

Abstract
The General Lagrangian Mean (GLM) theory uses a version of the averaged equations of fluid dynamics, designed to examine interactions between small-amplitude waves and mean flows. These equations are formulated in coordinates following the fluid’s average velocity and are often referred to as ‘pseudo-Lagrangian’. This paper focuses on the principles for deriving the GLM equations, using an inviscid, incompressible, homogeneous fluid as a demonstration case. Our exposition methodically differs from others and is aimed at the learners of this theory.
Keywords: fluid dynamics, GLM theory, inviscid incompressible fluid, Lagrangian displacements, mean flows, averaged equations.
1. Introduction
The General Lagrangian Mean (GLM) theory was introduced in 1978 by Andrews and McIntyre [1] to investigate wave-mean flow interaction in geophysical contexts, following Bretherton’s attempts [3]. An earlier version of this theory was studied in magnetohydrodynamics by Soward [7] in the context of a pseudo-Lagrangian interpretation of the ’nearly axisymmetric geodynamo theory’ [2]. These pioneering papers are difficult to follow in detail, whereas the detailed presentations of [4, 5] require professional efforts. The purpose of this paper is to present the underlying simple mathematical structure of GLM equations, with the aim of making the theory more accessible to a broad readership. In Section 2, we examine the mathematical structure of related transformations of Euler’s equations. Section 3 shows how these transformations connect to the GLM equations. The appendix presents the transformation of the continuity equation to illustrate the techniques involved.
2. What is the pseudo-Lagrangian description?
A fluid motion can be described by a link between the Lagrangian $\bm{X}$ and Eulerian $\bm{x}$ coordinates (and time $t$) of a fluid particle, $\bm{x}=\bm{x}(\bm{X},t),\ \bm{x}(\bm{X},0)=\bm{X}$, which we call an actual motion. Let us split $\bm{x}$ into two arbitrary parts $\bm{x}(\bm{X},t)\equiv\bar{\bm{x}}(\bm{X},t)+\bm{\xi}(\bm{X},t)$ and call $\bar{\bm{x}}$ the reference motion and $\bm{\xi}$ the secondary motion. Then, let us accept two one-to-one mappings $\bm{x}\Leftrightarrow\bm{X},\ \bar{\bm{x}}\Leftrightarrow\bm{X}$, which, by excluding $\bm{X}$, yields the one-to-one correspondence between the actual and reference motions,
$\bm{x}\Leftrightarrow\bar{\bm{x}},\quad\text{or}\quad\ \bm{x}=\bm{x}(\bar{\bm{x}},t),\ \bar{\bm{x}}=\bar{\bm{x}}(\bm{x},t).$ (2.1)
As a result, any function can be expressed in independent variables $(\bm{X},t)$, $(\bm{x},t)$, or $(\bar{\bm{x}},t)$. The description of motion $\bm{x}(\bar{\bm{x}},t)$ is called pseudo-Lagrangian, in contrast to Lagrangian one $\bm{x}(\bm{X},t)$. The two types of related velocities are
${\bm{u}}\equiv\partial{\bm{x}}/\partial t|_{\bm{X}},\qquad\bar{\bm{u}}\equiv\partial\bar{\bm{x}}/\partial t|_{\bm{X}},$ (2.2)
where the first is the actual velocity and the second is the reference velocity; the subscripts after the bars denote the variables kept constant. Physically, the smooth fields $\bar{\bm{x}}(\bm{X},t)$ and $\bar{\bm{u}}(\bm{X},t)$ represent arbitrary moving markers, chosen as a cloud of reference points. There are also three different expressions for the same ‘material’ derivative,
$D_{\bm{X}}\equiv{\frac{\partial}{\partial t}}\Big|_{\bm{X}}=\frac{\partial}{\partial t}\Big|_{\bm{x}}+{u}_{i}\frac{\partial}{\partial x_{i}}=\frac{\partial}{\partial t}\Big|_{\bar{\bm{x}}}+\bar{u}_{i}\frac{\partial}{\partial\bar{x}_{i}}\quad\text{or}\ \quad D_{\bm{X}}=D=\bar{D},$ (2.3)
where the summation convention over repeating vectorial subscripts is used. This symbolic equality among the three operators means that, when changing the independent variables, one should also automatically switch to the corresponding form of the operator. The next ansatz is the mutual split of both functions $\bm{x}$ and $\bar{\bm{x}}$ (2.1) into two addends,
$\bm{x}(\bar{\bm{x}},t)\equiv\bar{\bm{x}}+\bm{\xi}(\bar{\bm{x}},t)\Leftrightarrow\bar{\bm{x}}(\bm{x},t)\equiv\bm{x}+\bm{\eta}(\bm{x},t),$ (2.4)
From this point on, the independent variables used for $\bm{\xi}$ and $\bm{\eta}$ are:
$\bm{\xi}=\bm{\xi}(\bar{\bm{x}},t),\qquad\bm{\eta}=\bm{\eta}(\bm{x},t)=\bm{\eta}(\bar{\bm{x}}+\bm{\xi}(\bar{\bm{x}},t),t).$ (2.5)
The role of $\bm{\xi}$ and $\bm{\eta}$ is ‘to measure’ the deviation of the mapping $\bm{x}\Leftrightarrow\bar{\bm{x}}$ from the identity $\bm{x}=\bar{\bm{x}}$. An immediate outcome of (2.4) is:
$\bm{\eta}(\bm{x},t)=-\bm{\xi}(\bar{\bm{x}},t),\quad\text{or}\quad\bm{\eta}(\bar{\bm{x}}+\bm{\xi}(\bar{\bm{x}},t),t)=-\bm{\xi}(\bar{\bm{x}},t).$ (2.6)
Applying (2.3) to the first equality (2.4) yields
$\bar{D}\bm{\xi}=\bm{u}-\bar{\bm{u}}=-D\bm{\eta}.$ (2.7)
The Euler equations for an inviscid, incompressible, and homogeneous fluid are
$Du_{i}=-{\partial p}/{\partial x_{i}},\quad{\partial u_{k}}/{\partial x_{k}}=0,\quad D\equiv{\partial}/{\partial t}+{u_{k}}\,{\partial u_{i}}/{\partial x_{k}},\ \ i=1,2,3,\dots,$ (2.8)
where $\bm{x}=({x}_{1},{x}_{2},{x}_{3})$ are the actual Cartesian coordinates, $\bm{u}=(u_{1},u_{2},u_{3})$ is the actual velocity field, $p=p(\bm{x},t)$ is the pressure divided by the constant density. The straightforward change of independent variables $(\bm{x},t)\Rightarrow(\bar{\bm{x}},t)$, with the use of (2.1)-(2.7), transforms the first equation (2.8) into the form:
$\bar{D}^{2}\xi_{i}+\xi_{k,i}\bar{D}\bar{u}_{k}+\bar{D}(\bar{u}_{i}-\Pi_{i})+\bar{u}_{k,i}(\bar{u}_{k}-\Pi_{k})=-p^{*}_{,i},$ (2.9)
${\Pi}_{i}\equiv-{\xi}_{k,i}\bar{D}\xi_{k},\quad p^{*}\equiv p-u_{k}u_{k}/2.$
The shortcuts $(\cdot)_{,i}\equiv\partial/\partial\bar{x}_{i}$ and $(\cdot)_{,t}\equiv\partial/\partial t$ denote the partial derivatives; however, to avoid confusion, this notation is not used for the partial derivatives with respect to $\bm{x}$ and $\bm{X}$. The continuity equation (2.8), written in the variables $(\bar{\bm{x}},t)$, is (see the appendix)
${\partial u_{i}}/{\partial x_{i}}=[(\bar{D}\xi_{i})+\bar{u}_{i}]_{,k}(\delta_{ki}+\eta_{ki}(\bm{x},t))=\bar{u}_{i,i}+\bar{D}\xi_{i,i}+\eta_{ki}\bar{D}\xi_{i,k}=0,$ (2.10)
$\text{where}\quad\eta_{il}(\bm{x},t)\equiv\partial\bar{x}_{i}/\partial x_{l}=-J_{im}^{-1}\xi_{m,l},\ J_{im}\equiv\delta_{im}+\xi_{i,m},$
where $J_{im}^{-1}$ is a matrix, inverse to $J_{im}$. Expanding (2.10) for small $\xi_{i,k}$ and $\eta_{ik}$ of order $\varepsilon$ (A.5), yields
$\bar{u}_{i,i}=-\bar{D}\xi_{i,i}+\xi_{k,i}\bar{D}\xi_{i,k}+O(\varepsilon^{3}),$ (2.11)
which, in particular, shows that, contrary to the actual velocity $\bm{u}(\bm{x},t)$, the reference velocity field $\bar{\bm{u}}(\bar{\bm{x}},t)$ is ‘compressible’. Eqs.(2.9)-(2.11) represent a system of exact equations of fluid dynamics rewritten in independent variables $(\bar{\bm{x}},t)$. This system is underdetermined, as it contains seven unknown functions $\bar{\bm{u}}(\bar{\bm{x}},t)$, $\bm{\xi}(\bar{\bm{x}},t)$, and $p^{*}(\bm{x},t)$, but only four equations. This is natural, since six new unknown functions $\bar{\bm{u}}$ and $\bm{\xi}$ have been introduced in (2.1)-(2.4) to replace only three unknown functions $\bm{u}(\bm{x},t)$ in (2.8). Physically, since the transformation (2.1) contains an arbitrary function $\bar{\bm{x}}(\bm{x},t)$ that contains time, it represents a transfer from an inertial coordinate system $\bm{x}$ to a non-inertial one $\bar{\bm{x}}$. Therefore, we have the unusual appearance of the inertial (not real) forces in the transformed equations of motion. The approach considered can be called ‘pseudo-Lagrangian’ because instead of following material particles with coordinates $\bm{x}(\bm{X},t)$ (as in the Lagrangian approach), it follows arbitrarily chosen reference points $\bar{\bm{x}}(\bm{X},t)$. Here, a potential source of confusion is that all coordinate frames, $\bm{X}$, $\bm{x}$, and $\bar{\bm{x}}$, are fixed and Cartesian. However, the only coordinates $\bm{x}$ are inertial, while both $\bm{X}(\bm{x},t)$ and $\bar{\bm{x}}(\bm{x},t)$ are not.
3. Breakthrough to the major GLM equations
The GLM theory [1, 4, 5] exploits the mathematical freedom afforded by the previous transformations. The physical and mathematical meanings of $\bar{\bm{x}}$ are enforced by its expression through a set of actual motions $\bm{x}$ by an averaging operation, denoted $\langle\cdot\rangle$. The most popular is ensemble averaging,
$\langle\bm{x}\rangle=\bar{\bm{x}}\equiv\int{\bm{x}}(\bm{X},t,\alpha)f(\alpha)d\alpha,\ \text{with}\ \int f(\alpha)\,d\alpha=1,$ (3.1)
where $\alpha$ is an ensemble variable and $f(\alpha)\geq 0$ is a distribution function. The variable $\alpha$ in the text below is hidden. It is important that the averaging (3.1) commutes with all other mathematical operations used in this paper. The average operations in [1, 4, 5] vary. For establishing a relation with a wave motion and for the compatibility between (3.1) and (2.4), the secondary motion $\bm{\xi}$ must possess zero average and therefore is replaced by a more narrow class of functions $\bm{\xi}\Rightarrow\tilde{\bm{\xi}}=\tilde{\bm{\xi}}(\bar{\bm{x}},t)$, such that
$\langle\tilde{\bm{\xi}}\rangle\equiv 0,\quad\text{which entails}\quad\langle\bm{x}\rangle=\bar{\bm{x}},\quad\langle\bm{u}\rangle=\bar{\bm{u}},$ (3.2)
where the last two equalities are due to (2.4),(2.7),(3.1). From this stage, the field $\bar{\bm{x}}$ in (2.1) is not arbitrary and represents the mean motion, and $\tilde{\bm{\xi}}$ is the field of perturbations or the field of displacements. Applying the average operation reduces (2.9) to the equation,
$\bar{D}(\bar{u}_{i}-\bar{\Pi}_{i})+\bar{u}_{k,i}(\bar{u}_{k}-\bar{\Pi}_{k})=-\langle{p}^{*}\rangle_{,i}$
$\bar{\Pi}_{i}\equiv-\langle\tilde{\xi}_{k,i}\bar{D}\tilde{\xi}_{k}\rangle\equiv\bar{U}_{i}+\bar{A}_{ik}\bar{u}_{k};\quad\bar{U}_{i}\equiv-\langle\tilde{\xi}_{m,i}\tilde{\xi}_{m,t}\rangle,\ \bar{A}_{ik}\equiv-\langle\tilde{\xi}_{m,i}\tilde{\xi}_{m,k}\rangle,$
where $\bar{\bm{\Pi}}=(\bar{\Pi}_{1},\bar{\Pi}_{2},\bar{\Pi}_{3})$ is called the pseudomomentum vector. In turn, the average of the continuity equation (2.10) yields
$\bar{u}_{i,i}=-\langle(\bar{D}\tilde{\xi}_{i,k})\eta_{ki}\rangle,$ (3.4)
while its version for all $|\xi_{i,k}|\sim\varepsilon\ll 1$ is,
$\bar{u}_{i,i}=\langle\tilde{\xi}_{k,i}\bar{D}\tilde{\xi}_{i,k}\rangle+O(\varepsilon^{3}).$ (3.5)
The system of average equations (3.)-(3.5) conveniently connects the mean velocity $\bar{\bm{u}}$ and the displacement field $\tilde{\bm{\xi}}$. However, this system is undertermined (not closed).
4. How to solve the GLM equations?
One way of upgrading (3.)-(3.5) into a closed system is to obtain an independent evolution equation for the field $\tilde{\bm{\xi}}(\bar{\bm{x}},t)$. The most common approach [1, 4-6] is to consider small-amplitude waves and weak mean flow
$\tilde{\bm{\xi}}=O(\varepsilon),\quad p^{*}=O(\varepsilon),\quad\bar{\bm{u}}=O(\varepsilon^{2})\quad\text{with}\quad\varepsilon\ll 1.$ (4.6)
In this case, the linear evolution equations for $\tilde{\bm{\xi}}$ of order $O(\varepsilon)$ follow from (2.9)-(2.11),(3.1). As soon as it is solved, its solution is substituted in Eqs.(3.)-(3.5) (they are of order $O(\varepsilon^{2})$) for producing the function $\bar{\bm{u}}(\bar{\bm{x}},t)$. It represents the desired mean flow, linked to the linear wave $\tilde{\bm{\xi}}$. In summary, the closed system of equations in the approximation (4.6) is
$\tilde{\xi}_{i,tt}=-p^{*}_{,i},\quad\tilde{\xi}_{i,i}=0.\quad-\text{equations of order}\ O(\varepsilon),$ (4.7)
$\bar{D}(\bar{u}_{i}-\bar{\Pi}_{i})+\bar{u}_{k,i}(\bar{u}_{k}-\bar{\Pi}_{k})=-\langle{p}^{*}\rangle_{,i},\ \bar{u}_{i,i}=\langle\tilde{\xi}_{k,i}\bar{D}\tilde{\xi}_{i,k}\rangle-\text{of order}\ O(\varepsilon^{2}).$ (4.8)
The next step could be to develop this procedure by applying successive approximations; see [4]. Due to the simplest model (2.8) of fluid dynamics chosen, Eqn.(4.7) contains only potential solutions for $\tilde{\xi}_{i,tt}$, since any chosen flow (in its approximation) does not support wave-type motions. Several sample solutions with potential $\tilde{\bm{\xi}}$ can include combinations of point sources, vortices, or moving boundaries, which are beyond the scope of this paper, see [1, 4]. Grimshaw [7] provides a physically richer example of inhomogeneous (stratified) incompressible fluid flows that support internal waves. The final step of any GLM procedure is the return of the solutions $\bar{\bm{u}}(\bar{\bm{x}},t)$ and $\tilde{\bm{\xi}}(\bar{\bm{x}},t)$ to the actual (physical) independent variables $(\bm{x},t)$, which can be performed using (2.4)-(2.6). This operation can be physically sensible, for instance, when it changes the shapes of waves. However, for the approximation (4.6)-(4.8), one can use $\bar{\bm{x}}\simeq\bm{x}$, which also yields $\bm{\eta}\simeq-\tilde{\bm{\xi}}$. For a detailed exposition of the GLM theory with many examples, see [1, 4-7].
This paper was originally written and communicated to Professor A.D.D.Craik and Professor R.H.J.Grimshaw in 2009. Recently, it was edited using AI. The author acknowledges Professor H.K. Moffatt and Professor O. Bühler for interesting discussions.
5. Appendix: Transforming the continuity equation
To transform the continuity equation (2.8) into variables $(\bar{\bm{x}},t)$, we express $\bm{\eta}$ through $\bm{\xi}$ (2.4) using the identity
$J_{ik}J_{kl}^{-1}=\delta_{il},\quad\text{where}\quad J_{ik}\equiv{\partial x_{i}}/{\partial\bar{x}_{k}},\ J_{kl}^{-1}\equiv{\partial\bar{x}_{k}}/{\partial x_{l}},$ (A.1)
where $J_{kl}^{-1}$ is the inverse matrix to $J_{ik}$ and $\delta_{il}$ is the unit tensor. Eqn. (2.4) yields
$J_{ik}=\delta_{ik}+\xi_{i,k}(\bar{\bm{x}},t),\qquad J_{kl}^{-1}=\delta_{kl}+\eta_{kl}(\bm{x},t)=\delta_{kl}+\eta_{kl}(\bar{\bm{x}}+\bm{\xi},t),$ (A.2)
where $\eta_{kl}\equiv\partial\eta_{k}/\partial x_{l}$. Due to (A.1), the functions $\xi_{i,k}$ and $\eta_{kl}$ satisfy the equations,
$(\delta_{ik}+\xi_{i,k})(\delta_{kl}+\eta_{kl})=\delta_{il},\quad\text{or}\ \quad(\delta_{ik}+\xi_{i,k})\,\eta_{kl}=-\xi_{i,l},$ (A.3)
which represents a system of nine linear algebraic equations, with nine spatial derivatives, $\xi_{i,k}$, given, and nine unknowns, $\eta_{kl}$. It can be solved exactly and explicitly using standard linear algebra methods. One form of such a general solution is the following.
$\eta_{il}(\bm{x},t)\equiv-J^{-1}_{im}\xi_{m,l}=-\xi_{m,l}\{\delta_{im}+\xi_{i,m}\}^{-1}.$ (A.4)
In particular, this expression shows that the matrices $\eta_{ik}$ and $\xi_{i,k}$ commute, which is used to cancel specific terms in the transformed continuity equations. This commuting is because $\eta_{ik}$ represents a polynomial of $\xi_{i,k}$, which independent calculations can also confirm. For all $\xi_{i,k}$ and $\eta_{ik}$ having the same small amplitude $\varepsilon\ll 1$, it follows,
$\eta_{kl}=-\xi_{k,l}+\xi_{k,m}\xi_{m,l}+O(\varepsilon^{3}).$ (A.5)
The continuity equation in (2.8) takes the form,
${\partial u_{k}}/{\partial x_{k}}=[(\bar{D}\xi_{k})+\bar{u}_{k}]_{,l}J_{lk}^{-1}=0$ (A.6)
Using (2.7)(A.3)-(A.5) one can transform it to the required equations (2.10),(2.11).