### v2

# Discover the GLM theory on four pages

Abstract
The General Lagrangian Mean (GLM) theory uses a version of the averaged equations of fluid dynamics, designed to examine interactions between small-amplitude waves and mean flows. These equations are formulated in coordinates that follow the fluid’s average velocity and are often referred to as ‘pseudo-Lagrangian’ or ‘semi-Lagrangian’. This paper focuses on the principles for deriving the GLM equations, using an inviscid, incompressible, homogeneous fluid as a demonstration case. Our exposition methodically differs from others and is aimed at the learners of this theory.
Keywords: fluid flows, pseudo-Lagrangian description, GLM theory, inviscid incompressible fluid, Lagrangian displacements, mean flows, waves, averaged equations.
1. Introduction
The General Lagrangian Mean (GLM) theory was introduced in 1978 by Andrews and McIntyre [1] to investigate wave-mean flow interaction in geophysical contexts, following Bretherton’s attempts [3]. An earlier version of this theory was studied in magnetohydrodynamics by Soward [7] in the context of a pseudo-Lagrangian interpretation of the ’nearly axisymmetric geodynamo theory’ [2]. These pioneering papers are difficult to follow in detail, whereas reading monographs [4, 5] requires significant effort. Recent original publications are advanced and require passing the initial learning threshold; we avoid their review. The purpose of this paper is to expose the underlying mathematical origin of GLM equations with the aim of making the theory more accessible to a broad readership. Section 2 explains the meaning of the pseudo-Lagrangian description of fluid flow. Section 3 presents the Euler equations for an inviscid, incompressible, and homogeneous fluid in their pseudo-Lagrangian form. Section 4 shows how this form produces the averaged GLM equations. Section 5 contains a brief discussion on how to solve these equations. The appendix provides details on transforming the continuity equation into its GLM form to illustrate the techniques involved.
2. What is the ‘pseudo-Lagrangian’ description?
Let us introduce two smooth and invertible vector-functions in the three-dimensional space $\bm{X}=(X_{1}.X_{2},X_{3})$
$\bm{x}=\bm{x}(\bm{X},t),\qquad\bm{X}=\bm{X}(\bm{x},t),$ (2.1)
$\bar{\bm{x}}=\bar{\bm{x}}(\bm{X},t),\qquad\bm{X}=\bm{X}(\bar{\bm{x}},t)\,$ (2.2)
where (2.1) describes a fluid movement via the link between the Lagrangian $\bm{X}$ and Eulerian $\bm{x}$ coordinates (and time $t$) of the material particles, and (2.2) represents the motion of a cloud of markers that can be chosen arbitrarily, as convenient to a researcher. In the following, we call the former an actual motion, and the latter a reference motion. Two one-to-one mappings $\bm{x}\Leftrightarrow\bm{X},\ \bar{\bm{x}}\Leftrightarrow\bm{X}$ yield, by excluding $\bm{X}$, the one-to-one correspondence between the actual and reference motions:
$\bm{x}\Leftrightarrow\bar{\bm{x}},\quad\text{or}\quad\ \bm{x}=\bm{x}(\bar{\bm{x}},t),\ \bar{\bm{x}}=\bar{\bm{x}}(\bm{x},t).$ (2.3)
As a result, any function that describes the motion of a fluid can be expressed via independent variables $(\bm{X},t)$, $(\bm{x},t)$, or $(\bar{\bm{x}},t)$. The description of fluid motion $\bm{x}(\bar{\bm{x}},t)$ is called pseudo-Lagrangian, in contrast to the Lagrangian one $\bm{x}(\bm{X},t)$.
The two types of velocities are
${\bm{u}}\equiv\partial{\bm{x}}/\partial t|_{\bm{X}},\qquad\bar{\bm{u}}\equiv\partial\bar{\bm{x}}/\partial t|_{\bm{X}},$ (2.4)
where the first is the actual velocity and the second is the reference velocity; the subscripts after the bars denote the variables kept constant. Physically, the smooth fields $\bar{\bm{x}}(\bm{X},t)$ and $\bar{\bm{u}}(\bm{X},t)$ describe moving markers, chosen as reference points. There are three different expressions for the same ‘material’ derivative,
$D_{\bm{X}}\equiv{\frac{\partial}{\partial t}}\Big|_{\bm{X}}=\frac{\partial}{\partial t}\Big|_{\bm{x}}+{u}_{i}\frac{\partial}{\partial x_{i}}=\frac{\partial}{\partial t}\Big|_{\bar{\bm{x}}}+\bar{u}_{i}\frac{\partial}{\partial\bar{x}_{i}}\quad\text{or}\ \quad D_{\bm{X}}=D=\bar{D},$ (2.5)
where the summation convention over repeating subscripts is used. This symbolic equality of three operators means that changing the independent variables should automatically propagate to the corresponding forms of the operators.
For the transformation of the governing equation of motion to a pseudo-Lagrangian form, we use the example of an inviscid, incompressible, and homogeneous fluid,
$Du_{i}=-{\partial p}/{\partial x_{i}},\quad{\partial u_{k}}/{\partial x_{k}}=0,\quad D\equiv{\partial}/{\partial t}+{u_{k}}\,{\partial u_{i}}/{\partial x_{k}},\ \ i=1,2,3,\dots,$ (2.6)
where $\bm{x}=({x}_{1},{x}_{2},{x}_{3})$ are the actual Cartesian coordinates, $\bm{u}=(u_{1},u_{2},u_{3})$ is the actual velocity field, $p=p(\bm{x},t)$ is the pressure divided by the constant density. Then, the required pseudo-Lagrangian form of (2.6) is
$J_{ik}\bar{D}^{2}x_{i}=-p_{,k},\ J^{-1}_{ki}\partial u_{i}/\partial\bar{x}_{k}=0;\quad\bar{D}\equiv\partial/\partial t+\bar{u}_{k}\partial/\partial x_{k},\ J_{ik}\equiv{\partial x_{i}}/{\partial\bar{x}_{k}},$ (2.7)
where $J_{kl}^{-1}$ is the matrix inverse to $J_{ik}$ and the unknown functions are $\bm{x}(\bar{\bm{x}},t)$ and $\bar{\bm{u}}(\bar{\bm{x}},t)$. The latter represents the reference motion introduced by an arbitrary function in (2.2),(2.3). The shortcuts $(\cdot)_{,i}\equiv\partial/\partial\bar{x}_{i}$ and $(\cdot)_{,t}\equiv\partial/\partial t$ here and below denote the partial derivatives; however, to avoid confusion, this notation is not used for the partial derivatives with respect to $\bm{x}$ and $\bm{X}$.
3. Introducing Lagrangian displacements
The inconvenience of Eqs.(2.7) is its too general form, not adapted to the study of small perturbations of main flows. For making such adaptation, we split both functions (2.3) into two addends,
$\bm{x}(\bar{\bm{x}},t)\equiv\bar{\bm{x}}+\bm{\xi}(\bar{\bm{x}},t),\qquad\bar{\bm{x}}(\bm{x},t)\equiv\bm{x}+\bm{\eta}(\bm{x},t),$ (3.8)
where the roles of $\bm{\xi}$ and $\bm{\eta}$ are ‘to measure’ the deviation of the actual positions of fluid particles from the positions of markers or, mathematically, the deviation of the mapping $\bm{x}\Leftrightarrow\bar{\bm{x}}$ from the identity $\bm{x}=\bar{\bm{x}}$. The independent variables used in the following are
$\bm{\xi}=\bm{\xi}(\bar{\bm{x}},t),\qquad\bm{\eta}=\bm{\eta}(\bm{x},t)=\bm{\eta}(\bar{\bm{x}}+\bm{\xi}(\bar{\bm{x}},t),t),$ (3.9)
where $\bm{\xi}$ always has only this functional form, while $\bm{\eta}$ can be seen as a function of $(\bm{x},t)$ or $(\bar{\bm{x}},t)$. An immediate outcome of (3.8) is:
$\bm{\eta}(\bm{x},t)=-\bm{\xi}(\bar{\bm{x}},t),\quad\text{or}\quad\bm{\eta}(\bar{\bm{x}}+\bm{\xi}(\bar{\bm{x}},t),t)=-\bm{\xi}(\bar{\bm{x}},t).$ (3.10)
Applying (2.5) to (3.8) yields,
$\bar{D}\bm{\xi}=-D\bm{\eta}=\bm{u}-\bar{\bm{u}}.$ (3.11)
Its use along with (2.3)-(2.5) allows to transform the first equation (2.7), (2.6) into the form:
$\bar{D}^{2}\xi_{i}+\xi_{k,i}\bar{D}\bar{u}_{k}+\bar{D}(\bar{u}_{i}-\Pi_{i})+\bar{u}_{k,i}(\bar{u}_{k}-\Pi_{k})=-p^{*}_{,i},$ (3.12)
${\Pi}_{i}\equiv-{\xi}_{k,i}\bar{D}\xi_{k},\quad p^{*}\equiv p-u_{k}u_{k}/2,$
while the transformed continuity equation (2.7) is
$\bar{u}_{i,i}=-\bar{D}\xi_{i,i}-\eta_{ki}\bar{D}\xi_{i,k};\quad\text{where}\quad\eta_{il}(\bm{x},t)=-J_{im}^{-1}\xi_{m,l},\ J_{im}\equiv\delta_{im}+\xi_{i,m}.$ (3.13)
In particular, (3.13) shows that, contrary to the actual velocity $\bm{u}(\bm{x},t)$, the reference velocity field $\bar{\bm{u}}(\bar{\bm{x}},t)$ is ‘compressible’. Eqs.(3.12),(3.13) represent a system of exact equations of fluid dynamics rewritten in independent variables $(\bar{\bm{x}},t)$. This system is underdetermined, as it contains seven unknown functions $\bar{\bm{u}}(\bar{\bm{x}},t)$, $\bm{\xi}(\bar{\bm{x}},t)$, and $p^{*}(\bm{x},t)$, but only four equations. This is natural, since six new unknown functions $\bar{\bm{u}}$ and $\bm{\xi}$ have been introduced in (2.3)-(2.5) and (3.8) to replace only three unknown functions $\bm{u}(\bm{x},t)$ in (2.6).
Physically, since the transformation (2.3) contains an arbitrary function $\bar{\bm{x}}(\bm{x},t)$ depending on time, it represents a transfer from an inertial coordinate system $\bm{x}$ to a non-inertial one $\bar{\bm{x}}$. Therefore, we see the inertial (not real) forces that appear in the transformed equations of motion. The approach considered is called ‘pseudo-Lagrangian’ because, instead of considering material particles with Lagrangian coordinates $\bm{X}$ and related Eulerian coordinates $\bm{x}$, it uses arbitrarily chosen reference points $\bar{\bm{x}}$, also taken as Eulerian coordinates. A potential source of confusion is that all coordinate frames, $\bm{X}$, $\bm{x}$, and $\bar{\bm{x}}$, are fixed and Cartesian. However, the only coordinates $\bm{x}$ are inertial, while both $\bm{X}(\bm{x},t)$ and $\bar{\bm{x}}(\bm{x},t)$ are not.
4. Breakthrough to the GLM equations
The GLM theory [1, 4, 5] exploits the mathematical freedom afforded by the arbitrariness of the reference function $\bar{x}$. It enforces the physical and mathematical meanings of $\bar{\bm{x}}$ by expressing it as an averaged set of actual fluid motions $\bm{x}$. The average operations in [1, 4, 5] vary, while the most popular is ensemble averaging,
$\langle\bm{x}\rangle=\bar{\bm{x}}\equiv\int{\bm{x}}(\bm{X},t,\alpha)f(\alpha)d\alpha,\ \text{with}\ \int f(\alpha)\,d\alpha=1,$ (4.1)
where $\alpha$ is an ensemble variable and $f(\alpha)\geq 0$ is a distribution function. The variable $\alpha$ in the text below is hidden. It is important that the averaging (4.1) commutes with all other mathematical operations used in this paper. From this stage, the field $\bar{\bm{x}}$ in (2.3) is not arbitrary and represents the mean motion, and $\bm{\xi}$ is the field of perturbations or the field of displacements. For establishing a relation with a wave motion and for the compatibility between (4.1) and (3.8), the secondary motion $\bm{\xi}$ must possess zero average and therefore is replaced by a more narrow class of functions $\bm{\xi}\Rightarrow\tilde{\bm{\xi}}=\tilde{\bm{\xi}}(\bar{\bm{x}},t)$, such that
$\langle\tilde{\bm{\xi}}\rangle\equiv 0,\quad\text{which entails}\quad\langle\bm{x}\rangle=\bar{\bm{x}},\quad\langle\bm{u}\rangle=\bar{\bm{u}},$ (4.2)
where the last two equalities are due to (3.8),(3.11),(4.1).
Applying the average operation makes all linear terms in $\bm{\xi}$ disappear, which reduces (3.12) to the equation
$\bar{D}(\bar{u}_{i}-\bar{\Pi}_{i})+\bar{u}_{k,i}(\bar{u}_{k}-\bar{\Pi}_{k})=-\langle{p}^{*}\rangle_{,i}$ (4.3)
$\bar{\Pi}_{i}\equiv-\langle\tilde{\xi}_{k,i}\bar{D}\tilde{\xi}_{k}\rangle\equiv\bar{U}_{i}+\bar{A}_{ik}\bar{u}_{k};\quad\bar{U}_{i}\equiv-\langle\tilde{\xi}_{m,i}\tilde{\xi}_{m,t}\rangle,\ \bar{A}_{ik}\equiv-\langle\tilde{\xi}_{m,i}\tilde{\xi}_{m,k}\rangle,$
where $\bar{\bm{\Pi}}=(\bar{\Pi}_{1},\bar{\Pi}_{2},\bar{\Pi}_{3})$ is called the pseudomomentum vector. In turn, the average of the continuity equation (3.13) yields
$\bar{u}_{i,i}=-\langle(\bar{D}\tilde{\xi}_{i,k})\eta_{ki}\rangle,$ (4.4)
Equations (4.3),(4.4) conveniently connect the mean velocity $\bar{\bm{u}}$ and the displacement field $\tilde{\bm{\xi}}$. However, this system of equations is also underdetermined (not closed). The systems of equations (3.12),(3.13) and (4.3),(4.4) can be called ‘semi-Lagrangian’ because they contain both Lagrangian displacements $\bm{\xi}$ and Eulerian average velocity $\bar{\bm{u}}$.
5. How to solve the GLM equations?
One way of upgrading (4.3),(4.4) into a closed system is to obtain an independent evolution equation for the field $\tilde{\bm{\xi}}(\bar{\bm{x}},t)$. The most common approach [1, 4 - 6] is to consider small-amplitude waves and weak mean flow
$\tilde{\bm{\xi}}=O(\varepsilon),\quad p^{*}=O(\varepsilon),\quad\bar{\bm{u}}=O(\varepsilon^{2})\quad\text{with}\quad\varepsilon\ll 1.$ (5.5)
In this case, the linear evolution equations for $\tilde{\bm{\xi}}$ of order $O(\varepsilon)$ follow from (3.12)-(A.7),(4.1). As soon as it is solved, its solution is substituted in Eqs.(4.3),(4.4) (they are of order $O(\varepsilon^{2})$) to produce the function $\bar{\bm{u}}(\bar{\bm{x}},t)$. It represents the desired mean flow, linked to the linear wave $\tilde{\bm{\xi}}$. In summary, the closed system of equations in the approximation (5.5) is
$\tilde{\xi}_{i,tt}=-p^{*}_{,i},\quad\tilde{\xi}_{i,i}=0.\quad-\text{equations of order}\ O(\varepsilon),$ (5.6)
$\bar{D}(\bar{u}_{i}-\bar{\Pi}_{i})+\bar{u}_{k,i}(\bar{u}_{k}-\bar{\Pi}_{k})=-\langle{p}^{*}\rangle_{,i},\ \bar{u}_{i,i}=\langle\tilde{\xi}_{k,i}\bar{D}\tilde{\xi}_{i,k}\rangle-\text{of order}\ O(\varepsilon^{2}).$ (5.7)
The next step could be to develop this procedure by applying successive approximations; see [4]. Due to the simplest model (2.6) of fluid dynamics chosen, Eqn.(5.6) contains only potential solutions for $\tilde{\xi}_{i,tt}$, since any flow (in the chosen approximation) does not support wave-type motions. Several sample solutions with potential $\tilde{\bm{\xi}}$ can include combinations of point sources, vortices, or moving boundaries, which are beyond the scope of this paper, see [1, 4]. Grimshaw [7] provides a physically richer example of inhomogeneous (stratified) incompressible fluid flows that support internal waves.
The final step of the GLM procedure is the return of the solutions $\bar{\bm{u}}(\bar{\bm{x}},t)$ and $\tilde{\bm{\xi}}(\bar{\bm{x}},t)$ to the actual (physical) independent variables $(\bm{x},t)$, which can be performed using (3.8)-(3.10). This operation can be physically sensible, for instance, when it changes the shapes of waves. However, for the approximation (5.5)-(5.7), one can use $\bar{\bm{x}}\simeq\bm{x}$, which also yields $\bm{\eta}\simeq-\tilde{\bm{\xi}}$. For a detailed exposition of the variety of GLM equations and their solutions, see [1, 4 - 7].
Appendix: Continuity equation in the GLM theory
The continuity equation (2.6) with the use of (3.11) takes the form,
${\partial u_{k}}/{\partial x_{k}}=[(\bar{D}\xi_{k})+\bar{u}_{k}]_{,l}J_{lk}^{-1}=0,$ (A.1)
where we use the matrices
$J_{ik}\equiv{\partial x_{i}}/{\partial\bar{x}_{k}},\ J_{kl}^{-1}\equiv{\partial\bar{x}_{k}}/{\partial x_{l}},\quad\text{such that}\quad J_{ik}J_{kl}^{-1}=J_{ik}^{-1}J_{kl}=\delta_{il},$ (A.2)
where $J_{kl}^{-1}$ is the inverse matrix to $J_{ik}$ and $\delta_{il}$ is the unit tensor. Eqn. (3.8) yields
$J_{ik}=\delta_{ik}+\xi_{i,k}(\bar{\bm{x}},t),\qquad J_{kl}^{-1}=\delta_{kl}+\eta_{kl}(\bm{x},t)=\delta_{kl}+\eta_{kl}(\bar{\bm{x}}+\bm{\xi},t),$ (A.3)
where $\eta_{kl}\equiv\partial\eta_{k}/\partial x_{l}$. The expansion of both commuting equalities (A.2) shows that the functions $\xi_{i,k}$ and $\eta_{kl}$ satisfy the equations,
$\xi_{i,k}+\eta_{ik}+\xi_{i,n}\eta_{nk}=0,\qquad\xi_{i,k}+\eta_{ik}+\xi_{n,k}\eta_{in}=0,$ (A.4)
which means that the matrices $\xi_{i,k}$ and $\eta_{ik}$ also commute. The first equation (A.4) can be rewritten and solved as
$(\delta_{ik}+\xi_{i,k})\,\eta_{kl}=-\xi_{i,l},\quad\text{or}\ \quad\eta_{il}(\bm{x},t)\equiv-J^{-1}_{im}\xi_{m,l}=-\xi_{m,l}\{\delta_{im}+\xi_{i,m}\}^{-1}.$ (A.5)
Alternatively, Eqn.(A.5) can be solved exactly and explicitly using standard linear algebra methods. Indeed, it represents a system of nine linear algebraic equations with nine given coefficients $\xi_{i,k}$ and nine unknowns $\eta_{kl}$. Finally, the required GLM form of the continuity equation (3.13),(4.4) can be obtained using (A.1),(A.3),(A.4).
In an important special case where all $\xi_{i,k}$ and $\eta_{ik}$ have the same small amplitude $\varepsilon\ll 1$, it follows that
$\eta_{kl}=-\xi_{k,l}+\xi_{k,m}\xi_{m,l}+O(\varepsilon^{3}),$ (A.6)
where the reader can calculate any number of expansion terms. Expanding (3.13) for small $\xi_{i,k}$ of order $\varepsilon$, yields
$\bar{u}_{i,i}=-\bar{D}\xi_{i,i}+\xi_{k,i}\bar{D}\xi_{i,k}+O(\varepsilon^{3}),\qquad\bar{u}_{i,i}=\langle\tilde{\xi}_{k,i}\bar{D}\tilde{\xi}_{i,k}\rangle+O(\varepsilon^{3}),$ (A.7)
which confirms the second equation in (5.7).
This manuscript was initiated and written as the result of discussions with Professor A.D.D.Craik in 2009. The same year, it was communicated to Professor R.H.J.Grimshaw, and recently it was edited using artificial intelligence. The author acknowledges Professor H.K.Moffatt and Professor O.Bühler for recent interesting discussions.